OpenStudy (babyslapmafro):

Can someone please help explain the following... cot^4(pi-theta)=cot^4(theta) How are the above functions equal. What trig identity allows this to be true?

OpenStudy (babyslapmafro):

\[\cot ^{4}(\pi-\theta)=\cot ^{4}(\theta)\]

OpenStudy (anonymous):

Tangent and cotangent have a period of π.

OpenStudy (anonymous):

It's not really an identity (though, maybe it is depending on how you think about it), it's a property of the periodicity of cotangent. \[cot(\theta)=cot(\theta+n\pi)\]